# Convergent transformations of visual representation in brains and models

### v1

@@ A.2 Repetition-Matching Strategies for Inter-Subject RSA
A.3 Token-occurrence control for language–brain RSA
To quantify how much of the language–brain alignment reflects initial tokenization, we constructed RDMs from token-occurrence vectors. Each caption was encoded as a count vector—each entry recording the number of times a given token appears—and pairwise Pearson dissimilarities among these vectors yielded a tokenizer-based RDM. We applied this procedure to six vocabularies (BLOOMZ, Gemma 2, LLaMA 1/2/3; see Table 1), covering all tokenizers used in our language models, and then computed parcel-wise RSA against NSD fMRI responses using the symmetric HCP atlas. Parcel-wise alignments were nearly identical across tokenizers (pairwise Spearman’s $\rho$ $\leq 0.978$), so we report their average.
The tokenizer–brain alignment map closely matches that of full language models: it peaks in the LOTC hub and shows minimal or negative alignment in early visual parcels (Fig.A4A,D). Token-count geometry correlates strongly with each model’s first-layer alignment (Pearson’s $\rho=0.78$; Fig.A4B) and with maximal alignment across layers ($\rho=0.68$; Fig. A4C), indicating that token frequencies account for a substantial portion of the language–brain correspondence.
Repeating the analysis with a word tokenizer—splitting captions on whitespace rather than into subword units—again yielded pronounced LOTC alignment (Fig. A4E). These findings suggest that LOTC’s alignment to language models largely reflects sensitivity to discrete lexical tokens common across descriptions—functioning as a “bag-of-words” representation—rather than to compositional or syntactic structure53.
Figure A4: Tokenizer-level contributions to language–brain alignment. We derived RDMs from token-occurrence count vectors over each model’s vocabulary and computed parcel-wise RSA against NSD fMRI responses (averaged across the six tokenizers used in our language models). (A) Scatter of tokenizer-brain vs. inter-subject alignment, with peak values in the LOTC hub. (B) Token-count alignment against each model’s first-layer alignment. (C) Token-count alignment against each model’s maximal alignment across layers. (D) Cortical map of tokenizer-brain alignment, showing pronounced LOTC localization. (E) Alignment map for a whitespace-based word tokenizer, which likewise peaks in LOTC, consistent with a “bag-of-words” explanation rather than richer compositional structure.
A.4 Robustness to Similarity Metric (Spearman-RSA & CKA)
Recent work has cautioned that representational-similarity findings can depend not only on stimulus design but also on the choice of similarity metric—some measures probe local, pairwise geometry while others emphasize global, space-wide dependence 59. To test our results’ robustness to these different measures, we repeated key analyses using two alternatives to Pearson-RSA: (i) a rank-based Spearman-RSA, which preserves only the ordering of each stimulus pair’s dissimilarity, and (ii) an unbiased Centered Kernel Alignment (CKA) that quantifies global statistical dependence via an unbiased HSIC estimator 92.
Replacing Pearson with Spearman correlations yields a purely rank-based, nonparametric measure agnostic to absolute distance magnitudes. As shown in Supplementary Fig. A5A–C, the inter-subject map, the parcel-by-parcel scatter against model alignment, and the representational connectivity matrix all mirror our original Pearson-RSA results (parcel-wise $r=0.99$). This confirms that neither the linearity nor distributional assumptions of Pearson correlation drive our findings—our local pairwise geometry holds under a fully nonparametric test.
CKA measures dependence across entire representational spaces, emphasizing alignment along principal axes rather than isolated pairwise distances. We computed linear-kernel CKA (using the unbiased HSIC estimator 92) under the same shifted-repetition protocol. Supplementary Fig. A5D–F shows that CKA again recovers the three hubs but inflates LOTC alignment and attenuates early visual alignment in the parcel scatter, reflecting CKA’s sensitivity to low-dimensional, dominant components. Parcel-wise CKA and Pearson-RSA still correlate strongly (Pearson’s $r=0.91$, Spearman’s $\rho=0.93$), and the CKA connectivity matrix preserves the two-stream topology, despite CKA’s unsigned nature preventing a sign-based “dissimilarity” distinction.
Across both local (Spearman-RSA) and global (CKA) criteria, the identification of Early Visual, Ventral, and LOTC hubs and their two-stream connectivity remains unchanged. This convergence across similarity measures, spanning local versus global criteria, strengthens confidence that our observed hierarchies and pathways reflect genuine stimulus-driven geometry rather than artifacts of any single metric.
Figure A5: Comparison of alignment metrics. (A–C) Spearman-RSA. Replacing Pearson’s $r$ with Spearman’s rank correlation yields virtually identical maps: the inter-subject alignment (A), the parcel-wise scatter against vision- and language-model alignment (B), and the representational connectivity matrix (C) all match the Pearson-RSA findings, with parcel-wise strengths correlating at $r=0.99$. (D–F) Unbiased CKA. Centered Kernel Alignment, which captures global dependence via an unbiased HSIC estimator, again highlights Early Visual, Ventral, and LOTC hubs (D). In the parcel-wise comparison (E), CKA inflates LOTC alignment and attenuates early visual alignment—reflecting its sensitivity to low-dimensional, dominant components—yet CKA and Pearson-RSA remain strongly correlated ($r=0.91$). The CKA-based connectivity matrix (F) preserves the two-stream topology, though as a dependence measure it does not distinguish signed similarity.
A.5 Extended details of Shared Component Decomposition and Partial RSA Controls
Figure A6: Extended analysis of representational dimensions. (A) KMCCA projections onto the top two canonical axes for six representative parcels: V1, V4, VMV2, PHA2, MT, and TPOJ2. V1 shows no clear clustering; V4 begins to separate by content; Ventral and LOTC parcels replicate the hub-level semantic gradients. (B) Words ranked by loading on the first KMCCA component (positive loadings in blue, negative in red) for each hub. Early Visual words share low-level shape features; Ventral words span scene vs. object terms; LOTC words distinguish animate vs. inanimate labels. (C) Partial RSA boxplots for the three hubs, showing inter-subject alignment under no control (baseline), controlling for the scene-object gradient, or biological presence. LOTC alignment is markedly reduced by each control, whereas Early Visual and Ventral hubs are minimally affected.
This supplementary note details our use of Kernel Multi-view Canonical Correlation Analysis (KMCCA)33 to isolate the axes driving our RSA findings32. While RSA provides a single global similarity score, KMCCA decomposes this correspondence into a ranked set of orthogonal “representational channels,” revealing which latent axes contribute most strongly to the shared geometry.
We performed KMCCA using the Python package mvlearn (v0.5.0)90 with a correlation kernel and a regularization parameter of $\lambda=0.3$, using default parameters of implementation. For each cortical parcel (or hub), we formed a data matrix for each participant from the GLM-denoised single-trial $\beta$-estimates. We included only those trials for which each (image, repetition) pair was available for all eight participants. This procedure yielded a set of eight matrices ($X_{p,r}\in\mathbb{R}^{m\times v_{p,r}}$), one for each participant, all representing an identical sequence of stimulus trials. The KMCCA algorithm was then applied to this set of matrices, returning projections onto a common shared subspace. For visualization (Fig. A6A; Fig 5A-C), we projected each subject’s data onto the first two canonical axes and then averaged the projections for the repetitions of each unique image to produce a single, stable point per stimulus.
The parcel-level KMCCA projections (Supplementary Fig.A6A) confirmed that the organizational principles observed at the hub level were consistent within their constituent parcels. In the Early Visual Cortex, the V1 parcel showed no clear semantic organization in its first two dimensions, whereas the V4 parcel already exhibited emergent clustering by image content. Within the Ventral hub, representative parcels such as VMV2 and PHA2 reproduced the hub-level scene-to-object gradient and showed clear organization by semantic category. Similarly, within the LOTC hub, parcels like MT and TPOJ2 showed the same primary separation between biological and non-biological stimuli. Within these main clusters, stimuli were further organized by more fine-grained semantic categories, mirroring the main hub-level analysis (Fig.5C).
To interpret these KMCCA axes semantically, we projected the image captions onto the shared subspace learned for each hub. We first created a vocabulary from the union of all words present in the captions ($s=11,090$ unique words), encoding each caption as a binary word-occurrence vector (a “bag-of-words” representation). We then projected this caption matrix onto the first KMCCA component derived for each hub, which yielded a loading weight for each word in the vocabulary. By ranking these weights, we identified the words that contributed most positively and negatively to each dimension (Fig. A6B). This word-loading analysis confirmed our interpretation of each hub’s organizing principle: in Early Visual Cortex, top-weighted words were not semantically coherent but appeared to share coarse visual features; in the Ventral hub, positive tokens index scene contexts (e.g., ‘kitchen’, ‘bathroom’) while negative tokens name isolated objects (e.g., ‘plate’, ‘frisbee’); and in the LOTC hub, tokens cleanly separated animate from inanimate labels. This analysis provides a semantic interpretation for the organization of each cluster, which can be visually confirmed in the corresponding image atlases (See Supplementary Figs. D8-D10)
Finally, we quantified the impact of the identified dimensions on the overall inter-subject alignment using partial RSA. Denoting by $\rho(X,Y)$ the standard Pearson correlation use to compute the alignment between two RDMs, $Z$ a third RDM (e.g., one built from the first KMCCA component or from a binary “biological presence” vector), the partial RSA was computed as:
$\mathrm{pRSA}(X,Y\mid Z)=\frac{\rho(X,Y)\;-\;\rho(X,Z)\,\rho(Y,Z)}{\sqrt{(1-\rho(X,Z)^{2})\,(1-\rho(Y,Z)^{2})}}.$ (A1)
As shown in the main text (Fig. 5d-f), controlling for the continuous, data-driven KMCCA axis produced only minimal drops in the Early Visual and Ventral hubs, but a substantial reduction in the LOTC hub. A complementary analysis confirmed this result: controlling for the discrete categorical variables yielded a drop in alignment of a similar magnitude to that produced by the continuous KMCCCA components (Supplementary Fig. A6C).
A.6 Power-Law Attenuation Model of RSA Metrics
@@
In the main text, we compared several RSA measures—including within-subject (WS), inter-subject (IS-1: shift=1 and IS-0: unshifted), and vision-model-brain (VM)–that probed a common cortical geometry but were subject to different levels of measurement noise. While all metrics revealed a similar cortical alignment pattern, their magnitudes varied considerably. Parcel-wise scatter plots (e.g., Fig. 2E; Fig. A2) showed a systematic relationship between the metrics, which were tightly related by power-law curves. Although these curves differed in scale and curvature, their consistent form suggested that each metric captured the same latent similarity while imposing a unique, noise-dependent attenuation.
We formalise this observation by modelling the correlation measured for parcel r with metric j as a power-law transform of a latent correlation:
$\rho^{\mathrm{obs}}_{rj}=a_{j}\bigl{(}\rho^{\mathrm{latent}}_{r}\bigr{)}^{\,b_{j}}+\epsilon_{rj},$ (A2)
where
  • $\rho^{\mathrm{latent}}_{r}\in[0,1]$ is the metric-independent similarity common to all measures,
  • $a_{j}\in[0,1]$ is a linear attenuation factor for metric j, and
  • $b_{j}\geq 1$ captures signal-dependent (heteroscedastic) noise. A value of $b_{j}=1$ corresponds to simple homoscedastic attenuation, whereas $b_{j}>1$ creates a convex curve that suppresses weak latent correlations more strongly than strong ones. This constraint is motivated by our observation that the noisiest metrics (e.g., IS-shift) systematically underestimate weakly aligned parcels relative to cleaner metrics (e.g., WS), which is consistent with a noise regime that disproportionately affects weaker signals.
  • $\epsilon_{rj}$ is a residual term that absorbs any remaining parcel–metric–specific variability not captured by the deterministic power-law component.
We jointly fitted the parameters of Eq. A2 to four core metrics (WS, IS-0, IS-1, VM) using data from the symmetric-hemisphere ROIs. Parameter estimation was performed via the L-BFGS-B algorithm by minimizing a least-squares loss, with parameters bounded as follows: $a_{j}\in[0,1]$, $b_{j}\in[1,\infty)$, and $\rho^{\mathrm{true}}_{r}\in[0,1]$.
Figure A7A shows the model’s high goodness-of-fit by plotting the observed correlations against the model-fitted values for the four core metrics. In all cases, the model explained over 94% of the variance ($R^{2}>0.94$), confirming that the power-law attenuation model accurately predicts the trends in the observed data. Complementing this, Figure A7B visualizes the distinct attenuation profile for each RSA strategy, allowing for a direct observation of how each measured correlation is attenuated in comparison with the estimated latent correlation.
Because all metrics are modeled as functions of the same latent variable, $\rho^{\mathrm{true}}_{r}$, the relationship between any two metrics, j and j’, can be derived by eliminating this common term:
$y=\frac{a_{j^{\prime}}}{a_{j}^{\,b_{j^{\prime}}/b_{j}}}\;x^{\,b_{j^{\prime}}/b_{j}},$ (A3)
where $x=\rho^{\mathrm{obs}}_{rj}$ and $y=\rho^{\mathrm{obs}}_{rj^{\prime}}$. Figure A7C validates this prediction by overlaying the derived curve (Eq. A3) on the empirical parcel scatters. This confirmed that a single latent similarity, subject to metric-specific attenuation, explains all pairwise relations. For completeness, Figure A8 extends this validation to a set of 12 metrics, demonstrating the model’s power to unify all RSA strategies used in our analyses.
The resulting cortical map of the latent variable, $\rho^{\mathrm{latent}}_{r}$, is displayed in Figure A7D. This map represents a noise-corrected estimate of similarity that unifies the four core RSA measures. It reveals a three-hub pattern of high alignment observed in the main text, with peak correlations ($\rho^{\mathrm{latent}}_{r}\sim 0.50$) in the EVC, Ventral stream, and LOTC. Furthermore, the model shows that prefrontal areas (FEF and pFC) are more prominent after correcting for the attenuation present in the individual raw measurements. The fact that these hubs align neatly with those highlighted by individual metrics confirms that our model (Eq. A2) successfully isolates a common signal while factoring out metric-specific noise.
Because Eq. (A2) is scale-free, absolute values of $\hat{\rho}^{\mathrm{latent}}$ and $a_{j}$ are defined only up to a common factor. We fixed no anchor during optimisation, yet repeated fits converged to the same scale, and all qualitative findings (relative parcel ranking, cross-metric relations) are invariant under rescaling.
The power-law attenuation model offers a compact, quantitative account of how measurement differences degrades RSA scores across diverse metrics. Its success supports the view that all metrics probe a shared latent representational geometry, differing only in the reliability with which each metric samples that geometry.
@@
Figure A7: (A) Goodness-of-fit. Parcel-wise observed RSA values (y-axis) are plotted against the corresponding predictions from the joint power-law model (Eq. A2; x-axis). Colours denote the different metrics (WS: within-subject; IS-1: inter-subject shifted (shift=1); IS-0: inter-subject unshifted; VM: vision model). The legend provides the explained variance ($R^{2}$) for each metric. (B) Correlation attenuation curves. The estimated latent correlation ($\hat{\rho}^{\mathrm{latent}}$, x-axis) is plotted against the observed correlation for each metric (y-axis). Coloured lines show the fitted metric-specific attenuation curves ($y=a_{j}\,x^{b_{j}}$). (C) Pairwise metric comparisons. Scatter plots for six representative metric pairs (grey points) are shown. Each plot includes an independent power-law fit to the empirical data (solid red line, with 95% CI) and the prediction derived from the joint model (dashed black line, Eq. A3), demonstrating high correspondence. (D) Latent similarity map. The cortical surface map displays the estimated latent correlation values, $\hat{\rho}^{\mathrm{latent}}_{r}$ (HCP-MMP1 atlas, symmetric parcels). The model reveals major hubs of high latent similarity in early visual cortex, ventral temporal cortex, and LOTC, with pFC becoming more prominent after accounting for metric-specific adjustment. Together, these panels demonstrate that a single latent geometry, subject to metric-specific power-law attenuation, effectively explains the relationships among all measured RSA metrics.

### v2

@@ Acknowledgments
Supplementary Note 2: Parcel size and the reliability of IS-RSA estimates
HCP–MMP parcels vary substantially in surface area and voxel count, raising the concern that IS-RSA estimates might be systematically less stable in smaller parcels due to reduced signal averaging. To evaluate this, we quantified the dispersion of IS-RSA estimates under two resampling schemes that isolate distinct sources of variability: changes in stimulus sampling and trial-level measurement noise.
To assess stability with respect to stimulus sampling, we recomputed IS-RSA across 1,000 random half-splits of the image set (500 unique images per split, including all repetitions; Supplementary Fig. 14a-d). For each parcel, we computed the coefficient of variation (CV; standard deviation of the distribution divided by the mean) as a scale-normalized index of estimation stability.
To isolate trial-level measurement variability while holding stimulus content fixed, we performed 1,000 iterations in which we randomly selected a single repetition per image (from the 2–3 available) and recomputed IS-RSA (Supplementary Fig. 14e-h). Dispersion was again summarized per parcel using the CV.
Across both analyses, parcels with similar mean IS-RSA exhibited comparable dispersion despite large differences in voxel count (Supplementary Fig.14a-b, e-f). Critically, CV showed no systematic dependence on parcel size (Supplementary Fig.14c,g). Instead, dispersion decreased monotonically with the magnitude of IS-RSA (Supplementary Fig. 14d,h), indicating that parcels with stronger shared geometry yield more stable IS-RSA estimates regardless of their size. These results suggest that, under the NSD preprocessing regime, the reliability of IS-RSA estimates is primarily governed by the strength of the shared representational signal rather than by the number of voxels available for averaging.
Figure 14: Quantifying parcel reliability via resampling. (A–D) Stimulus generalization analysis. We computed IS-RSA on 1,000 random half-splits of the stimulus set (selecting 500 stimuli in each subsample). (A,B) Distribution of IS-RSA scores for a representative large parcel (V1) and small parcel (VMV2). Dashed vertical lines indicate the IS-RSA estimate from the full dataset. (C) IS-RSA coefficient of variation (CV; std/mean) plotted against parcel voxel count. (D) CV plotted against mean IS-RSA alignment. (E–H) Measurement reliability analysis based on randomly selecting one repetition per image across 1,000 iterations. (E,F) Corresponding distributions for V1 and VMV2. (G) CV versus voxel count. (H) CV versus mean IS-RSA alignment. In both analyses, the dispersion shows no dependence on voxel count but decreases as a function of the IS-RSA magnitude.
Supplementary Note 3: Hemispheric Asymmetry and the Validity of Symmetric Analyses
Figure 15: Hemispheric comparison of representational alignment. (A-C) Comparison of representational alignment (RSA, Pearson’s $r$; $N=8$, NSD) computed independently for the left (x-axis) and right (y-axis) hemispheres: (A) inter-subject alignment, (B) vision-model alignment, and (C) language-model alignment. Each point is a cortical parcel, colored by macro-anatomical group. The diagonal line indicates equal alignment. Squares denote parcels with a statistically significant hemispheric difference (paired $t$-test, $p<0.05$, FDR-corrected). (D) Boxplots of alignment within principal hubs. (E) Correlation between cross-hemispheric and bilateral inter-subject alignment across parcels. Significant hemisphere effects are marked.
We quantified hemispheric asymmetry in representational geometry by computing alignment measures separately for the left and right hemispheres using the HCP-MMP1.0 atlas 34. For each cortical parcel, we assessed inter-subject alignment, vision-model alignment, and language-model alignment using the same NSD responses as in the main analyses (Fig. 15).
Across all modalities, a largely symmetric pattern was observed (Fig. 15a-c). The rank-order of parcels was highly consistent between hemispheres (Spearman’s $\rho$: inter-subject =0.93=0.93, vision =0.92=0.92, language =0.82=0.82). However, specific asymmetries emerged: early visual cortex (V1–V4) consistently showed stronger alignment in the left hemisphere, while the ventral and LOTC hubs (e.g., hMT+, LO, TPOJ) showed stronger alignment in the right hemisphere (Fig. 15d). These lateralization trends were stable across inter-subject and model-based analyses.
To determine whether these asymmetries reflected fundamental differences in representational geometry or merely differences in signal strength, we compared cross-hemispheric RSA (e.g., subject $p$’s left vs subject $q$’s right) to bilateral RSA. These measures were near-perfectly correlated across parcels ($r=0.99$, Fig. 15e), with no parcel deviating substantially from the regression line. This demonstrates that combining hemispheres increases overall alignment strength (likely via noise averaging) without distorting the underlying representational geometry; no region exhibited a lateralized geometry that was absent in the contralateral hemisphere.
Together, these results validate our strategy of pooling hemispheres for the main analyses: aggregation maximizes statistical power for multivariate comparisons without introducing representational artifacts, while still allowing for targeted lateralization analyses where relevant.
Supplementary Note 4: Power-Law Attenuation Model of RSA Metrics
@@
In the main text, we compared several RSA measures–including within-subject (WS), inter-subject (IS-1: shift=1 and IS-0: unshifted), and vision-model-brain (VM)–that probed a common cortical geometry but were subject to different levels of measurement noise. While all metrics revealed a similar cortical alignment pattern, their absolute magnitudes varied considerably. Parcel-wise scatter plots (e.g., Fig. 2e; Fig. 12) showed a systematic relationship between the metrics, which were tightly related by power-law curves. Although these curves differed in scale and curvature, their consistent functional form suggested that each metric captured the same latent similarity structure while imposing a unique, noise-dependent attenuation.
@@
Figure 16: Power-law attenuation model. (A) Goodness-of-fit. Parcel-wise observed RSA values (y-axis) plotted against predictions from the joint power-law model (Eq. 1; x-axis). Colors denote metrics (WS: within-subject; IS-1: inter-subject shifted; IS-0: inter-subject unshifted; VM: vision model). Legend indicates explained variance ($R^{2}$). (B) Attenuation curves. Estimated latent correlation ($\hat{\rho}^{\mathrm{latent}}$, x-axis) vs. observed correlation (y-axis). Colored lines show the fitted metric-specific attenuation curves ($y=a_{j}x^{b_{j}}$). (C) Pairwise comparisons. Scatter plots for six representative metric pairs (gray points). Each plot includes an independent power-law fit (solid red line, 95% CI) and the prediction from the joint model (dashed black line, Eq. 2), demonstrating high correspondence. (D) Latent similarity map. Cortical surface map of the estimated latent correlation values, $\hat{\rho}^{\mathrm{latent}}_{r}$ (HCP-MMP1 atlas, symmetric parcels). The model recovers the three major hubs (EVC, Ventral, LOTC) and suggests distinct signal in prefrontal regions after accounting for metric-specific attenuation.
We formalize this observation by modeling the correlation measured for parcel $r$ with metric $j$ as a power-law transform of a latent correlation:
$\rho^{\mathrm{obs}}_{rj}=a_{j}\bigl(\rho^{\mathrm{latent}}_{r}\bigr)^{\,b_{j}}+\epsilon_{rj},$ (1)
where:
  • $\rho^{\mathrm{latent}}_{r}\in[0,1]$ is the metric-independent similarity common to all measures;
  • $a_{j}\in[0,1]$ is a linear attenuation factor for metric $j$;
  • $b_{j}\geq 1$ captures non-linear signal compression. A value of $b_{j}=1$ corresponds to simple linear attenuation, whereas $b_{j}>1$ creates a convex curve that suppresses weak latent correlations more strongly than strong ones. This constraint is motivated by the observation that noisier metrics (e.g., IS-shift) systematically underestimate weakly aligned parcels relative to cleaner metrics (e.g., WS), consistent with a noise regime that disproportionately obscures weaker signals;
  • $\epsilon_{rj}$ is a residual term absorbing parcel–metric–specific variability not captured by the deterministic power-law component.
We jointly fitted the parameters of Eq. 1 to four core metrics (WS, IS-0, IS-1, VM) using data from the symmetric-hemisphere ROIs. Parameter estimation was performed via the L-BFGS-B algorithm by minimizing the sum of squared residuals, with bounds: $a_{j}\in[0,1]$, $b_{j}\in[1,\infty)$, and $\rho^{\mathrm{latent}}_{r}\in[0,1]$.
Figure 16A demonstrates the model’s high goodness-of-fit. In all cases, the model explained over 94% of the variance ($R^{2}>0.94$), confirming that the power-law formulation accurately captures the relationship between observed data and the latent construct. Figure 16B visualizes the distinct attenuation profile for each RSA strategy, illustrating how each metric compresses the estimated latent correlation.
Because all metrics are modeled as functions of the same latent variable, $\rho^{\mathrm{latent}}_{r}$, the relationship between any two metrics, $j$ and $j^{\prime}$, can be derived analytically by eliminating this common term:
$y=\frac{a_{j^{\prime}}}{a_{j}^{\,b_{j^{\prime}}/b_{j}}}\;x^{\,b_{j^{\prime}}/b_{j}},$ (2)
where $x=\rho^{\mathrm{obs}}_{rj}$ and $y=\rho^{\mathrm{obs}}_{rj^{\prime}}$. Figure 16C validates this prediction by overlaying the derived curve (Eq. 2) on the empirical parcel scatters. The tight alignment confirms that a single latent similarity, subject to metric-specific attenuation, explains the pairwise relationships. For completeness, Figure 17 extends this validation to a set of 12 metrics, demonstrating the model’s ability to unify all RSA strategies used in our analyses.
The resulting cortical map of the latent variable, $\hat{\rho}^{\mathrm{latent}}_{r}$, is displayed in Figure 16D. This map represents a attenuation-corrected estimate of similarity that unifies the core RSA measures. It recovers the three-hub pattern observed in the main text, with peak correlations ($\rho^{\mathrm{latent}}_{r}\sim 0.50$) in the EVC, Ventral stream, and LOTC. Furthermore, the model suggests that frontal areas (e.g., FEF and PFC) become more prominent after correcting for the attenuation present in raw measurements. The fact that these hubs align with those highlighted by individual metrics confirms that Eq. 1 successfully isolates a common signal while factoring out metric-specific noise.
Note that because Eq. 1 is scale-free, the absolute values of $\hat{\rho}^{\mathrm{latent}}$ and $a_{j}$ are defined only up to a common scaling factor. We fixed no anchor during optimization, yet repeated fits converged to the same scale, and all qualitative findings (relative parcel ranking, cross-metric relations) are invariant under rescaling.
The power-law attenuation model offers a compact, quantitative account of how measurement differences degrade RSA scores. Its success supports the view that all metrics probe a shared latent representational geometry, differing only in the fidelity with which each metric samples that geometry.